if father is 4 time older from daughter, and 6 years back he was 10 time older what is his age?
Answer by Alan3354(69443)

if father is 4 time older from daughter, and 6 years back he was 10 time older what is his age?
---------------
f = father's age, d = daughter's
---
4 times older = 5 times as old
now:
f = 5d
6 yrs ago:
f-10 = 11*(d-10)
f-10 = 11d - 110
f = 11d - 100
Sub for f
5d = 11d - 100
-6d = -100
d = 16 2/3 years
f = 83 1/3 years
-------------------
The answers are not integers because you're expected to think that "4 times older" means 4 times as old.
It does not.
Using %ages, if something is 100% older --> 2x as old.
1 time older = 2 times as old.
----------
Tell the teacher the author of the textbook should find a different line of work.